The table summarizes results from pedestrian deaths that were caused by automobile accidents.

Pedestrian Deaths
Driver
Intoxicated?
Pedestrian Intoxicated?
Yes No
Yes 55 77
No 233 611

If one of the pedestrian deaths is randomly selected, find the probability that the pedestrian was intoxicated. (Round your answer to 4 decimal places.)
Probability =

B.

Giving a test to a group of students, the grades and gender are summarized below. Round your answers to 4 decimal places.

Grades and Gender
A B C Total
Male 16 17 13 46
Female 11 20 19 50
Total 27 37 32 96


If one student is chosen at random,

a. Find the probability that the student got an A:  

b. Find the probability that the student was male AND got an "A":  

c. Find the probability that the student was female OR got a "C":  

d. If one student is chosen at random, find the probability that the student got an 'B' GIVEN they are male:
Answer #1
Pedestrian Deaths
Driver
Intoxicated?
Pedestrian Intoxicated?
Yes No
Yes 55 77
No 233 611

the probability that the pedestrian was intoxicated. =  

the probability that the pedestrian was intoxicated. = 0.2951

ans 2 )

Grades and Gender
A B C Total
Male 16 17 13 46
Female 11 20 19 50
Total 27 37 32 96

a. the probability that the student got an A = 27/96 =0.2813

b the probability that the student was male AND got an "A" = 16/ 46 =0.3478

c. There are 46 males total and 32 people who got a C  So there are 46+32 = 78 people who either are male or got a B

We need to subtract off the 13 males who got C . i.e 78 -13 = 65 now there are 96 student total

P(male OR got a C) = (number who are male or got a C or both)/(total no. of students )
P(male OR got a C) = 65/96
P(male OR got a C) = 0.6771

d)  one student is chosen at random, the probability that the student got an 'B' GIVEN they are male = 16 / 27 =0.5926